OpenStudy (anonymous):

I'll fan and medal // Use a transformation to solve the equation. f – 12 = –35 A. f = –47 B. f = –23 C. f = 23 D. f = 47 I think the answer is -23 ,but i'm not sure

OpenStudy (misty1212):

i think so too

OpenStudy (misty1212):

\[f-12=-35\\ f=-35+12=-23\]

OpenStudy (anonymous):

Thank you Misty1212!